June 13, 1999 OWL shift summary

Richard Kessler (rkessler@fnal.gov)
Sun, 13 Jun 1999 07:54:16 -0500 (CDT)

Shift Summary for OWL shift, 6/13/99

On Shift:RSK, RP

Hours of beam delivered at all intensities: 1.5
Hours of beam delivered at nominal intensity: 1.5
Hours of E832 data (special runs not included): 1
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: ktev2e disk filled up

========================================================================
Runs ended this shift:
Run 13553 - 1.8M events - BEAM_MUON
Run 13555 - 3.9M events - BEAM_MUON

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
We finally got beam at 5am and started with muons [run 13553].
ktev2e disk filled up and run was aborted. Expert was paged
and lead us through the cache-delete procedure. Two previous
muon caches were backed up but not deleted which caused the
problem.

After an hour of muons [runs 13553-555]we switched to hadrons with
the accidental disk file writing turned back on. We removed the
nominal tapes and put in dummies in case it hangs with no EOF mark.
Just as we finished intializing a quench occurred.

When there was no beam, we sent a pulser into the accid 90 trigger
and wrote those events to disk with BEAM_832 config. This worked
fine for three consecutive runs. So we suspect the problem may only
occur when all tapes and files are used ... i.e., we may really
need 800 GeV protons to solve this software problem ???

The monitor-run-ntuple has been revived.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

========================================================================
Follow up Notes on Previous Shift

CsI survived the night with no more casualties.
One channel has a gain drift of -7 %.

We noticed that the mean momentum in the two beams for
the good BEAM_832 run 13542 differs by 0.3%.

Concerning the 30% charged 2pi mode excess, the Z and mass
distributions (w/cuts) are identical to those in 1997.
See logbook distributions. So still no clue about this.

========================================================================
Items to Follow up on Next Shift

accid 90 disk file is turned ON.
The goal is to re-produce the failure and note carefully
exactly what happens:
The turn off disk-writing in disk832.nml
(remove file 6), put back nominal tapes and take data.

** END ***